OpenStudy (anonymous):

Solve the system by substitution. 2x+y=-11 3x-4y=11 a:(3,5) b:(-5,-3) c:(-3,-5) d:(5,3)

OpenStudy (anonymous):

Solve for x on the first equation. Then *substitute* that value into the second equation.

OpenStudy (anonymous):

What do you get when you solve for x on the first equation?

OpenStudy (anonymous):

x=-y/2-11/2

OpenStudy (anonymous):

@muntoo

OpenStudy (anonymous):

Now plug the value of x you just got into the second equation, 3x-4y=11.

OpenStudy (anonymous):

After I plugged it in and solved for y I got -3 @muntoo

OpenStudy (anonymous):

\[\large{ 3x - 4y = 11 \\ \ \\ 3({-y\over 2}-{11\over 2}) - 4y = 11 \\ \ \\\ 3({-y-11\over 2}) - 4y = 11 \\ \ \\ 3(-y-11) - 8y = 22 \\ -3y-33 - 8y = 22 \\ -11y = 55 \\ y = -5 }\] Now plug that value of y back into one of the equations (any one) to find x.